Find the difference between the sum of the first positive 100 multiples of 3 and the sum of the first 100 positive even integers.

Answer by mananth(16946)

multiples of 3
3,6,9,.........
n=100
a=3
d=3
sn=n/2[2a+(n-1)d)]
=>100/2[6+99*3]
=>50*303
=> 15150
---
even integers
a=2
d=2
n=100
Sn=n/2[2a+(n-1)d]
=50[4+99*2]
=50(202)
=10100
Find the difference